# AI-Driven Intrusion Detection Systems (IDS) on the ROAD dataset: A Comparative Analysis for automotive Controller Area Network (CAN)

### v1

@@ 4.4. TAN
The last approach for intrusion detection that we adopted in this study relies on a transformer-based architecture developed by Nguyen et al. (10141599). Differently from the original transformer architecture (vaswani2023attention), the TAN architecture only uses the encoder, the most essential component for classification, thus resembling the highly influential ViT model produced by Google (dosovitskiy2021image). Furthermore, the model has been tested for transfer learning, which aims to address a significant challenge in in-vehicle intrusion detection. This method enables the application of a model trained on a large amount of data from a specific vehicle to a different vehicle for which less data is available.
@@ 5.1. Comprehensive Framework: Experiment Setup, Data Preprocessing, and Performance Evaluation Metrics
It’s worth noting that the ROAD dataset is highly imbalanced, as it can be seen on Table 1, and some classes need to be oversampled. For this reason we employed the SMOTE algorithm (Chawla_2002) on the training set to increase the number of samples of each attack to a fixed value of 100,000, therefore ensuring an equal distribution between them, that otherwise could have not been detected from any metric produced by our models after building the sequences. The In-Vehicle Network Intrusion Detection Challenge dataset has been oversampled in the same way. The resulting data can be directly fed to the traditional machine learning models and the LSTM-based model, but it needs to undergo further processing in order to create the frames needed by the DCNN and the sequences needed by the TAN, built as described in Section 4, in Figure 2 and Figure 3.
@@ 5.2.1. LCCDE and ML approaches
The performance of the LCCDE model, consisting of three gradient boosting classifiers (LightGBM, CatBoost and XGBoost), has been reproduced on the HCRL Car-Hacking dataset and we obtained very good and constant results on the In-Vehicle Network Intrusion Detection Challenge dataset. However, the model has not performed as well on the ROAD dataset, despite having analyzed both the ID and data field, with an average F1 score of 0.3212 and standard deviation of 0.1723. It is worth noting that this model could not detect at all two of the five masquerade attacks. A similar situation involves Random Forests which on average outperform slightly LightGBM, but had worse results than LCCDE. Furthermore, the LCCDE model was the one that required the most computation time due to the original implementation and impossibility of parallelization, requiring about 32 hours and 40 minutes of running time for the classification task on the ROAD dataset, while LightGBM has only taken about 6 minutes, XGBoost has taken about 10 minutes and CatBoost has taken about one hour and 27 minutes, which means that the vast majority of the computation time is due to the LCCDE algorithm. The code for the LCCDE model was made available by the authors (lccde-code).
@@ 5.2.3. DCNN
We have been able to reproduce the performance shown on the HCRL Car-Hacking dataset, which was built by the same authors of this model (SONG2020100198). The results were not as stable on the In-Vehicle Network Intrusion Detection Challenge dataset, having a higher average standard deviation of 0.05 on the F1 score metric, while the model has seen worse performance and dramatic fluctuations on the ROAD dataset. This has probably been caused by the fact that this dataset contains attacks acting also on the data field and they are much more difficult to detect because of their stealthiness, described on the Section 3.1. Considering this, the DCNN has performed fairly well, sometimes producing less false positives than the TAN model, with a notable exception for the fuzzing attack. This attack is easily identifiable by the payload, always set on the maximum value, but this model only uses the ID field, that in this case has cycling IDs. The cycling pattern is hardly detectable from the input frame containing 29 consecutive IDs, which explains the unreliability of the model.

### v2

@@ 4.4. TAN
The last approach for intrusion detection that we adopted in this study relies on a transformer-based architecture developed by Nguyen et al. (Nguyen et al., 2023). Differently from the original transformer architecture (Vaswani et al., 2023), the TAN architecture only uses the encoder, the most essential component for classification, thus resembling the highly influential ViT model produced by Google (Dosovitskiy et al., 2021). Furthermore, the model has been tested for transfer learning, which aims to address a significant challenge in in-vehicle intrusion detection. This method enables the application of a model trained on a large amount of data from a specific vehicle to a different vehicle for which less data is available.
@@
[…]
[The picture shows the sequence creation process for the TAN model]Similarly as for the DCNN, sequences with step=1 are created and if they contain an attack frame’s ID, shown in red, then they are labelled as attack (1), 0 otherwise.
@@ 5.1. Comprehensive Framework: Experiment Setup, Data Preprocessing, and Performance Evaluation Metrics
It’s worth noting that the ROAD dataset is highly imbalanced, as it can be seen on Table 1, and some classes need to be oversampled. For this reason we employed the SMOTE algorithm (Chawla et al., 2002) on the training set to increase the number of samples of each attack to a fixed value of 100,000, therefore ensuring an equal distribution between them, that otherwise could have not been detected from any metric produced by our models after building the sequences. The In-Vehicle Network Intrusion Detection Challenge dataset has been oversampled in the same way. The resulting data can be directly fed to the traditional machine learning models and the LSTM-based model, but it needs to undergo further processing in order to create the frames needed by the DCNN and the sequences needed by the TAN, built as described in Section 4, in Figure 2 and Figure 3.
@@ 5.2.1. LCCDE and ML approaches
The performance of the LCCDE model, consisting of three gradient boosting classifiers (LightGBM, CatBoost and XGBoost), has been reproduced on the HCRL Car-Hacking dataset and we obtained very good and constant results on the In-Vehicle Network Intrusion Detection Challenge dataset. However, the model has not performed as well on the ROAD dataset, despite having analyzed both the ID and data field, with an average F1 score of 0.3212 and standard deviation of 0.1723. It is worth noting that this model could not detect at all two of the five masquerade attacks. A similar situation involves Random Forests which on average outperform slightly LightGBM, but had worse results than LCCDE. Furthermore, the LCCDE model was the one that required the most computation time due to the original implementation and impossibility of parallelization, requiring about 32 hours and 40 minutes of running time for the classification task on the ROAD dataset, while LightGBM has only taken about 6 minutes, XGBoost has taken about 10 minutes and CatBoost has taken about one hour and 27 minutes, which means that the vast majority of the computation time is due to the LCCDE algorithm. The code for the LCCDE model was made available by the authors (Lab, 2021).
@@ 5.2.3. DCNN
We have been able to reproduce the performance shown on the HCRL Car-Hacking dataset, which was built by the same authors of this model (Song et al., 2020). The results were not as stable on the In-Vehicle Network Intrusion Detection Challenge dataset, having a higher average standard deviation of 0.05 on the F1 score metric, while the model has seen worse performance and dramatic fluctuations on the ROAD dataset. This has probably been caused by the fact that this dataset contains attacks acting also on the data field and they are much more difficult to detect because of their stealthiness, described on the Section 3.1. Considering this, the DCNN has performed fairly well, sometimes producing less false positives than the TAN model, with a notable exception for the fuzzing attack. This attack is easily identifiable by the payload, always set on the maximum value, but this model only uses the ID field, that in this case has cycling IDs. The cycling pattern is hardly detectable from the input frame containing 29 consecutive IDs, which explains the unreliability of the model.